# Late Jets, Early Sparks: Illuminating the Pre-Maximum Bumps in Superluminous Supernovae

### v1

@@ 2.1 Simulation setup
We carry out a 3D relativistic hydrodynamic simulation of a jet launched into an expanding stellar envelope, using the code pluto (Mignone et al., 2007). We follow the jet propagation in the ejecta, its activity after breakout, and after it shuts off. We post-process the simulation output semi-analytically to produce estimates of the post-breakout emission. As the simulations are dimensionless, we can calibrate the output with dimensionless mass scale $m_{s}$ and timescale (length-scale) $t_{s}$.
The jet operates for $\,{t_{\rm eng}}=4460\,t_{s}\,\,{{\rm s}}$ with a total (two-sided) luminosity, $\,{L_{\rm j}}=2\times 10^{47}\,m_{s}/t_{s}\,{\rm erg\,\,{{\rm s}}^{-1}}$. We consider as fiducial values $\{t_{s}=80,m_{s}=1\}$, corresponding to jet duration $\,{t_{\rm eng}}\approx 4.1\,\,{{\rm d}}$, and luminosity $L=2.5\times 10^{45}\,{\rm erg}\,\,{{\rm s}}^{-1}$, typical of the engine properties needed to power SLSNe-I (Fig. 1; see also Metzger et al. 2015, their Fig. 1). These jet properties are illustrated as an orange square in Fig. 1, while left-right arrows indicate the change in engine properties obtained by varying $t_{s}$ ($L_{\rm eng}\propto t_{\rm eng}^{-1}$) that would result in the same hydrodynamic evolution (for fixed ejecta properties $m_{s}$). This degeneracy running approximately parallel to the SLSNe population suggests that the hydrodynamic evolution predicted by our single simulation could apply to a wide range of SLSNe engines.
Building on the results of Gottlieb et al. (2022a), we set the jets at the inner boundary to be precessing about the spin-axis of the black hole, with an inclination angle of $0.05$ rad and a period of $t_{s}\,\,{{\rm s}}$. The initial jet Lorentz factor is $\Gamma_{0}=10$ with a maximum Lorentz factor [^2] We note that the precise ultra-relativistic Lorentz factor of the jet does not affect the cocoon, which maintains $\Gamma\lesssim 3$ at all times (Gottlieb et al., 2021). of 20 (equipartition between initial thermal and kinetic energy), and its opening angle is $\theta_{0}=0.1$ rad. The jet runs into a homologously expanding $M_{\rm ej}=10\,m_{s}\,{M_{\odot}}$ ejecta with a front velocity of $\,{v_{\rm ej}}=10^{4}\,\,{{\rm km}}\,\,{{\rm s}}^{-1}$ and initial position $\,{r_{\rm ej,0}}=7\times 10^{6}\,t_{s}\,\,{{\rm km}}$. The ejecta initial comoving mass density profile is:
@@
We use an ideal gas equation of state with a polytropic index of 4/3, as appropriate for radiation-dominated gas. This applies at all relevant times for our calculation as we focus on the outflow while radiation is coupled to the gas. The 3D Cartesian grid is divided into 3 patches on the $\hat{x}$- and $\hat{y}$-axes: The inner patch is at the innermost $2\times 10^{5}t_{s}\,\,{{\rm km}}$ with 100 uniformly distributed cells. The outer patches are stretched from $\lvert 2\times 10^{5}\,t_{s}\,\,{{\rm km}}\rvert$ to $\lvert 3\times 10^{8}\,t_{s}\,\,{{\rm km}}\rvert$ with 350 logarithmically distributed cells. We divide the $\hat{z}$-axis (jet axis) into 4 patches: starting from $z_{0}=2\times 10^{5}\,t_{s}\,\,{{\rm km}}$ to $3\times 10^{6}\,t_{s}\,\,{{\rm km}}$ with 300 uniformly distributed cells, followed by 3 logarithmically spaced patches with 600 cells to $3\times 10^{7}\,t_{s}\,\,{{\rm km}}$, 350 cells to $9.5\times 10^{7}\,t_{s}\,\,{{\rm km}}$, and 480 cells to $6\times 10^{8}\,t_{s}\,\,{{\rm km}}$. The jet is injected at the lower boundary, $z_{0}$, with a cylindrical injection radius of $r_{0}=z_{0}\theta_{0}$. The integration is performed with a Harten-Lax-van Leer (HLL) solver, Runge-Kutta time stepping, and piecewise parabolic interpolation.
@@ 2.2 Emission calculation
where $h$ and $k_{B}$ are the Planck and Boltzmann constants, respectively, and $\nu_{\rm max}(\bar{t})={\rm max}\left[\nu L_{\nu}(\bar{t})\right]$.
@@ 3.1 NUV/optical
Figure 2: 3D rendering of the logarithmic comoving mass density at the end of the jet activity, $t=\,{t_{\rm eng}}$, when the ejecta front is at $\,{r_{\rm ej}}\approx 3.5\times 10^{14}\,\,{{\rm km}}$. As the jet escapes from the high-density expanding spherical ejecta (yellow), it generates a stratified structure of the lower-density cocoon (red). The interaction of the jet and the cocoon with the expanding ejecta close to the breakout radius generates bright NUV/optical emission over a $\sim$ week timescale.
Figure 3: 2D $\hat{x}-\hat{z}$-plane cuts at various times of the azimuthally-averaged logarithmic comoving thermal energy density $3p/c^{2}$ maps. The white and black contours represent the photosphere and trapping radius, respectively.
Figure 4: (a): Bolometric luminosity peaks soon after breakout ($t_{b}$), and then declines as $\sim t^{-2}$ (dashed black line). The dashed magenta lines illustrate the rise of the main SLSN-I peak. (b): Color temperature of the radiating gas. (c): Absolute (left vertical axis) and apparent (right vertical axis) magnitude at various bands ($f_{\rm NUV}=1.2\times 10^{15}\,{\rm Hz}$, $f_{g}=6.3\times 10^{14}\,{\rm Hz}$, $f_{r}=4.8\times 10^{14}\,{\rm Hz}$, $f_{y}=3.0\times 10^{14}\,{\rm Hz}$, $f_{K_{s}}=1.4\times 10^{14}\,{\rm Hz}$) demonstrate that the peak magnitude aligns with observations of SLSN early bumps. (d): Luminosity-weighted average dimensionless velocity of the radiating gas at the trapping radius demonstrates that the emitting gas is always subrelativistic. The dimensionless front ejecta velocity is shown for comparison in red.
The head of the weak jet propagates slowly inside the expanding ejecta with an average velocity in the ejecta frame of $v_{h}-\,{v_{\rm ej}}\approx 0.01c$. This implies that all the jet energy is practically deposited into the cocoon for as long as the jet head remains within the ejecta. For our choice of parameters: $t_{s}=80$ and $m_{s}=1$, the two-sided jet with a luminosity of $\,{L_{\rm j}}=2.5\times 10^{45}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ breaks out from the $M_{\rm ej}=10\,\,{M_{\odot}}$ ejecta after $t_{b}\approx 2.2\,\,{{\rm d}}$, indicating that the cocoon energy upon breakout is $L_{c}\approx\,{L_{\rm j}}t_{b}\approx 5\times 10^{50}\,\,{{\rm erg}}$. Figure 2 illustrates a 3D rendering of the logarithmic comoving mass density, showcasing the interaction of the jet-cocoon outflow (red) with the homologously expanding ejecta (yellow). This interaction generates emission near the optically thin breakout radius and shapes the large-scale stratified structure observed beyond the confined expanding ejecta.
@@
(1) The first emission phase begins upon breakout and continues for as long as the jet engine is active at $t_{b}<t<t_{\rm eng}$. Fig. 3(a) depicts the breakout of the collimated jet and its hot, pressurized cocoon from the expanding ejecta. Upon breakout, photons above the trapping radius escape from the top layers of the cocoon, producing the first light through shock breakout emission. At this time, the emitting layers are mildly relativistic, as indicated by the proximity of the trapping radius to the photosphere. Shortly after the shock breakout signal, the gas from deeper layers starts to contribute to the emission with $\,{\beta_{\rm emit}}\approx 0.2$ [Fig. 4(d)], and the luminosity begins to fall off [Fig. 4(a)]. The high temperature after breakout of $T_{c}\approx 2\times 10^{4}\,{\rm K}$ [Fig. 4(b)] results in a peak in the NUV [Fig. 4(c)].
After the breakout, the jet head accelerates to ultra-relativistic velocities, and the mildly relativistic cocoon expands in the angular direction to $\theta\sim(\Gamma\beta)^{-1}$. In typical GRBs, the breakout takes place within $\sim 100\,\,{{\rm s}}$, implying that the photons generated in the jet-cocoon interaction are still trapped well within the outflow. Here, Figure 3(b), which depicts the jet-cocoon system at the time of jet engine shut-off at $t_{\rm eng}=4.1\,\,{{\rm d}}$, suggests that the jet is optically thin, and the late breakout results in a trapping radius located just outside of the jet-cocoon interface (JCI; Gottlieb et al., 2021) and within the ejecta outer radius. Consequently, the emission originates from the continuous heating in the JCI, as indicated by the high thermal energy along the increasing trapping radius in the angular direction away from the jet axis. Over time, the drop in the temperature [Fig. 4(b)] results in a transition to softer emission [Fig. 4(c)].
@@
The high temperatures around the time of breakout cause the emission to first peak in the NUV and then in the optical bands, as the latter overlaps with the blackbody peak while the jet is active. As the luminosity declines more rapidly than the temperature at $t\gtrsim t_{\rm eng}$, the NIR $K_{s}$ band emission rises since it is located on the Rayleigh-Jeans portion of the spectrum. Consequently, the decrease in the bolometric luminosity following jet shutdown also reduces the optical emission.
(3) Comparing Figures 3(c) and 3(d), it is evident that the trapping radius shrinks in the ejecta frame as the ejecta becomes optically thin. The final phase of the emission will occur once the cocoon crosses the entire unshocked ejecta, and the entire ejecta becomes optically thin. The timing of this transition depends on the opacity, which may decrease with temperature, depending on the composition of the supernova ejecta. In particular, if the ejecta is primarily comprised of O-Ne-Mg nuclei, the electron scattering and line-opacity drops substantially once $T\approx 6\times 10^{3}\,{\rm K}$, at $t\approx 5\,\,{{\rm d}}$ due to recombination effects (e.g., Kleiser et al. 2018). However, if the line opacity of Fe-group nuclei (e.g., Fe-Co-Ni) dominates, the opacity will remain relatively constant across this temperature range, causing the optically thin transition to occur somewhat later. We note that the relevant opacity for the emission is determined below the trapping radius, where the temperature is $T_{t}=T_{c}\,{\beta_{\rm emit}}^{-0.25}\approx 2T_{c}$, implying that $T_{t}\gtrsim 6\times 10^{3}\,{\rm K}$ at $t\leq 16\,{\rm d}$. This suggests that the numerical calculation remains valid at the relevant times.
Ultimately, the rise of the main supernova emission, powered primarily by (a more isotropic, non-jet-related) source of ejecta heating by the same central engine (e.g., the absorption of X-rays and gamma-rays from the magnetar nebula; Vurm & Metzger 2021), will take over the optical light curve. The dashed magenta curves in the bolometric luminosity panel illustrate the rise of the SLSN-I emission towards the main peak at $L_{\rm bol}\approx 10^{44}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ at $t\approx 50\,\,{{\rm d}}$ (e.g., Quimby et al., 2011; Gal-Yam, 2012). We present both a slow rise, $L_{\rm bol}\sim t$, and a fast rise, $L_{\rm bol}\sim t^{2}$.
@@ 3.2 Radio
Ultimately, the jet will decelerate due to its interaction with the dense CSM, powering an “orphan” radio afterglow synchrotron emission and an optical flash from the reverse shock (Sari & Piran, 1999). However, the latter is likely to be outshone by the thermal cocoon emission associated with the early bump. While the lack of orphan afterglow detections in blind surveys (Ho et al., 2018, see however Perley et al. 2024 for a recent potential detection), along with the potential for high CSM densities around some SLSNe-I (e.g., Margutti et al., 2023) that would amplify the signal, may argue against the generic presence of jets in these systems, the large uncertainties prevent us from reaching definitive constraints.
Figure 5: Top-hat jet afterglow light curves at $\nu=6\,{\rm GHz}$ for observers at $\theta_{\rm obs}=0.5\,{\rm rad}$ (solid lines) and on-axis observers (dashed lines) at $z=0.25$, assuming jet Lorentz factor $\Gamma=100$ and opening angle $\theta_{j}=0.1\,{\rm rad}$, microphysics parameters: electron equipartition parameter $\epsilon_{e}=0.1$, magnetic equipartition parameter $\epsilon_{B}=10^{-4}$, and electron energy power-law index $p=2.2$, and wind velocity $v_{w}=1000\,\,{{\rm km}}\,\,{{\rm s}}^{-1}$. The black dots denote the most constraining upper limits measured for SLSN-I Gaia2016apd (Coppejans et al., 2018), placing the source at $z=0.25$.
Figure 5 presents the semi-analytic thin-shell approximation solution for the radio synchrotron emission of our model at various assumed progenitor wind mass-loss rates, for both off-axis (solid lines) and on-axis (dashed lines) viewing angles. We assume a characteristic redshift of $z=0.25$ for SLSNe-I (e.g., Gomez et al., 2024), standard jet properties with $\Gamma=100$ and $\theta_{j}=0.1\,{\rm rad}$, and microphysics parameters (detailed in the caption) inferred from GW170817 afterglow and VLBI images (Mooley et al., 2018). If the mass-loss rates are $M_{w}\gtrsim 10^{-5}\,{\rm yr}^{-1}$, the emission is self-absorbed during the first $\sim 1$ week, leading to an initial rise in the light curves for all observers.
@@
Beyond the general uncertainties in the afterglow and jet parameters, the strong dependence of the flux on the viewing angle (Nakar et al., 2002) and the wind mass-loss rate complicates robust predictions that can support or refute our model. The most stringent upper limits on SLSNe-I were obtained for Gaia2016apd (Coppejans et al., 2018) at $z=0.1018$ (Yan et al., 2017). The black dots represent these upper limits (adjusted for $z=0.25$), potentially ruling out some scenarios with high mass-loss rates and small viewing angles for this source, but not excluding off-axis observers and lower mass-loss rates within the considered parameters.

### v2

@@ 2.1 Simulation setup
We carry out a 3D relativistic hydrodynamic simulation of a jet launched into an expanding stellar envelope, using the code pluto (Mignone et al., 2007). We follow the jet propagation in the ejecta, its activity after breakout, and the period after it shuts off. We post-process the simulation output semi-analytically to produce estimates of the post-breakout emission. As the simulations are dimensionless, we can calibrate the output with dimensionless mass scale $m_{s}$ and timescale (length scale) $t_{s}$.
The jet operates for $\,{t_{\rm eng}}=4460\,t_{s}\,\,{{\rm s}}$ with a total (two-sided) luminosity, $\,{L_{\rm j}}=2\times 10^{47}\,m_{s}/t_{s}\,{\rm erg\,\,{{\rm s}}^{-1}}$. We consider as fiducial values $\{t_{s}=200,m_{s}=3\}$, corresponding to jet duration $\,{t_{\rm eng}}=10.3\,\,{{\rm d}}$, and luminosity $L_{\rm j}=3\times 10^{45}\,{\rm erg}\,\,{{\rm s}}^{-1}$, typical of the engine properties needed to power Type I SLSNe (Fig. 1; see also Metzger et al. 2015, their Fig. 1). These jet properties are illustrated as an orange square in Fig. 1, while bidirectional arrows indicate the change in engine properties obtained by varying $t_{s}$ ($L_{\rm eng}\propto t_{\rm eng}^{-1}$) that would result in the same hydrodynamic evolution (for fixed ejecta properties $m_{s}$). This degeneracy running approximately parallel to the SLSN population suggests that the hydrodynamic evolution predicted by our single simulation could apply to a wide range of SLSN engines. For the early bump to be observable, the jet breakout must occur early enough to avoid being outshone by the primary SN emission. Eq. 1 demonstrates that, for typical jet and ejecta properties, the breakout timescale is indeed expected to be significantly shorter than the peak of the main SLSN.
Building on the results of Gottlieb et al. (2022a), we set the jets at the inner boundary to be precessing about the spin axis of the black hole, with an inclination angle of $0.05$ rad and a period of $t_{s}\,\,{{\rm s}}$. The initial jet Lorentz factor is $\Gamma_{0}=10$ with a maximum Lorentz factor [^2] We note that the precise ultra-relativistic Lorentz factor of the jet does not affect the cocoon, which maintains $\Gamma\lesssim 3$ at all times (Gottlieb et al., 2021). of 20 (equipartition between initial thermal and kinetic energy), and its opening angle is $\theta_{0}=0.1$ rad. The jet runs into a homologously expanding $M_{\rm ej}=7\,m_{s}\,{M_{\odot}}$ ejecta with a front velocity of $\,{v_{\rm ej}}=10^{4}\,\,{{\rm km}}\,\,{{\rm s}}^{-1}$ and initial position $\,{r_{\rm ej,0}}=7\times 10^{6}\,t_{s}\,\,{{\rm km}}$. The ejecta initial comoving mass density profile is:
@@
We use an ideal gas equation of state with a polytropic index of 4/3, as appropriate for radiation-dominated gas. This applies at all relevant times for our calculation as we focus on the outflow while radiation is coupled to the gas. The 3D Cartesian grid is divided into three patches on the $\hat{x}$- and $\hat{y}$-axes: The inner patch is at the innermost $2\times 10^{5}\,t_{s}\,\,{{\rm km}}$ with 100 uniformly distributed cells. The outer patches are stretched from $\lvert 2\times 10^{5}\,t_{s}\,\,{{\rm km}}\rvert$ to $\lvert 3\times 10^{8}\,t_{s}\,\,{{\rm km}}\rvert$ with 350 logarithmically distributed cells. We divide the $\hat{z}$-axis (jet axis) into four patches: starting from $z_{0}=2\times 10^{5}\,t_{s}\,\,{{\rm km}}$ to $3\times 10^{6}\,t_{s}\,\,{{\rm km}}$ with 300 uniformly distributed cells, followed by three logarithmically spaced patches with 600 cells to $3\times 10^{7}\,t_{s}\,\,{{\rm km}}$, 350 cells to $9.5\times 10^{7}\,t_{s}\,\,{{\rm km}}$, and 480 cells to $6\times 10^{8}\,t_{s}\,\,{{\rm km}}$. The jet is injected at the lower boundary, $z_{0}$, with a cylindrical injection radius of $r_{0}=z_{0}\theta_{0}$. The integration is performed with a Harten-Lax-van Leer (HLL) solver, Runge-Kutta time stepping, and piecewise parabolic interpolation.
@@ 2.2 Emission calculation
where the integration is performed over one hemisphere, as the diffusion time through the equator exceeds the dynamical time ($\tau\beta>1$), restricting observers near the axis to observing only one jet.
@@
where $h$ and $k_{B}$ are the Planck and Boltzmann constants, respectively, and $\nu_{\rm max}(\bar{t})={\rm arg\,max}\left[L_{\nu}(\bar{t})\right]$.
@@ 3.1 NUV/Optical
Figure 2: 3D rendering of the logarithmic comoving mass density at the end of the jet activity, $t=\,{t_{\rm eng}}$, when the ejecta front is at $\,{r_{\rm ej}}\approx 8.5\times 10^{14}\,\,{{\rm km}}$. As the jet escapes from the high-density expanding spherical ejecta (yellow), it generates a stratified structure of the lower-density cocoon (red). The interaction of the jet and the cocoon with the expanding ejecta close to the breakout radius generates bright NUV/Optical emission over a $\sim$ week timescale.
Figure 3: 2D $\hat{x}-\hat{z}$-plane cuts at various times of the azimuthally-averaged logarithmic comoving thermal energy density, $e=3p$ maps. The white and black contours represent the photosphere and trapping radius, respectively.
Figure 4: (a): Bolometric luminosity peaks soon after breakout ($t_{b}$), and then declines. The dashed magenta lines illustrate the rise of the main Type I SLSN peak. (b): Color temperature of the radiating gas. (c): Absolute (left vertical axis) and apparent (right vertical axis) magnitude at various bands ($f_{\rm NUV}=1.2\times 10^{15}\,{\rm Hz}$, $f_{u}=8.6\times 10^{14}\,{\rm Hz}$, $f_{g}=6.3\times 10^{14}\,{\rm Hz}$, $f_{r}=4.8\times 10^{14}\,{\rm Hz}$, $f_{y}=3.0\times 10^{14}\,{\rm Hz}$) demonstrate that the peak magnitude aligns with observations of SLSN early bumps. (d): Luminosity-weighted average dimensionless velocity of the radiating gas at the trapping radius demonstrates that the emitting gas is always subrelativistic. The dimensionless front ejecta velocity is shown in red for comparison.
The head of the weak jet propagates slowly inside the expanding ejecta with an average velocity in the ejecta frame of $v_{h}-\,{v_{\rm ej}}\approx 0.01c$. This implies that all the jet energy is practically deposited into the cocoon for as long as the jet head remains within the ejecta. For our choice of parameters: $t_{s}=200$ and $m_{s}=3$, the two-sided jet with a luminosity of $\,{L_{\rm j}}=3\times 10^{45}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ breaks out from the $M_{\rm ej}\approx 20\,\,{M_{\odot}}$ ejecta after $t_{b}\approx 5.5\,\,{{\rm d}}$, indicating that the cocoon energy upon breakout is $L_{c}\approx\,{L_{\rm j}}t_{b}\approx 1.7\times 10^{51}\,\,{{\rm erg}}$. Figure 2 illustrates a 3D rendering of the logarithmic comoving mass density, showcasing the interaction of the jet-cocoon outflow (red) with the homologously expanding ejecta (yellow). This interaction generates emission near the optically thin breakout radius and shapes the large-scale stratified structure observed beyond the confined expanding ejecta.
@@
(1) The first emission phase begins upon breakout and continues for as long as the jet engine is active at $t_{b}<t<t_{\rm eng}$. Fig. 3(a) depicts the breakout of the collimated jet and its hot, pressurized cocoon from the expanding ejecta. Upon breakout, photons above the trapping radius escape from the top layers of the cocoon, producing the first light through shock breakout emission. At this time, the emitting layers are mildly relativistic, as indicated by the proximity of the trapping radius to the photosphere. Shortly after the shock breakout signal, the gas from deeper layers starts to contribute to the emission with $\,{\beta_{\rm emit}}\approx 0.2$ [Fig. 4(d)], and the luminosity begins to fall off [Fig. 4(a)]. The high temperature immediately after breakout of $T_{c}\approx 7\times 10^{4}\,{\rm K}$ [Fig. 4(b)] results in dominant ultraviolet emission [Fig. 4(c)].
After the breakout, the jet head accelerates to ultra-relativistic velocities, and the mildly relativistic cocoon expands in the angular direction to $\theta\sim(\Gamma\beta)^{-1}$. In typical GRBs, the breakout takes place within $\sim 100\,\,{{\rm s}}$, implying that the photons generated in the jet-cocoon interaction are still trapped well within the outflow. Here, Figure 3(b), which depicts the jet-cocoon system at the time of jet engine shutoff at $t_{\rm eng}=10.3\,\,{{\rm d}}$, suggests that the jet is optically thin, and the late breakout results in a trapping radius located just outside of the jet-cocoon interface (JCI; Gottlieb et al., 2021) and within the ejecta outer radius. Consequently, the emission originates from the continuous heating in the JCI, as indicated by the high thermal energy along the increasing trapping radius in the angular direction away from the jet axis. Over time, the drop in the temperature [Fig. 4(b)] results in a transition to NUV emission which peaks at $t\approx t_{\rm eng}$ [Fig. 4(c)].
@@
The high temperatures at jet activity cause the emission to first peak in the ultraviolet, followed by a peak in the optical bands after the jet engine is shut off. As the luminosity declines more rapidly than the temperature at $t\gtrsim t_{\rm eng}$, the near-infrared $y$band emission rises since it is located on the Rayleigh-Jeans portion of the spectrum. Consequently, the decrease in the bolometric luminosity following jet shutdown also reduces the optical emission.
(3) Comparing the panels in Fig. 3, it is evident that the trapping radius shrinks in the ejecta frame, and the entire unshocked ejecta has become optically thin in Fig. 3(d). The final phase of the emission will occur once the cocoon becomes optically thin. The timing of this transition depends on the opacity, which may decrease with temperature, depending on the composition of the SN ejecta. In particular, if the ejecta is primarily composed of O-Ne-Mg nuclei, the electron scattering and line-opacity drops substantially once $T\approx 6\times 10^{3}\,{\rm K}$, at $t\approx 5\,\,{{\rm d}}$ due to recombination effects (e.g., Kleiser et al. 2018). However, if the line opacity of Fe-group nuclei (e.g., Fe-Co-Ni) dominates, the opacity will remain relatively constant across this temperature range, causing the optically thin transition to occur somewhat later.
Ultimately, the rise of the main SN emission, powered primarily by a (more isotropic, non-jet-related) source of ejecta heating by the same central engine (e.g., the absorption of X-rays and gamma-rays from the magnetar nebula; Vurm & Metzger 2021), will take over the optical light curve. The dashed magenta curves in the bolometric luminosity panel illustrate the rise of the Type I SLSN emission towards the main peak at $L_{\rm bol}\approx 10^{44}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ at $t\approx 50\,\,{{\rm d}}$ (e.g., Quimby et al., 2011; Gal-Yam, 2012). We present both a slow rise, $L_{\rm bol}\sim t$, and a fast rise, $L_{\rm bol}\sim t^{2}$.
@@ 3.2 Radio
Ultimately, the jet will decelerate due to its interaction with the dense CSM, powering an “orphan” radio afterglow synchrotron emission and an optical flash from the reverse shock (Sari & Piran, 1999). However, the latter is likely to be outshone by the thermal cocoon emission associated with the early bump. While the lack of orphan afterglow detections in blind surveys (Ho et al., 2018, see however Perley et al. 2024 for a recent potential detection), along with the potential for high CSM densities around some Type I SLSNe (e.g., Margutti et al., 2023) that would amplify the signal, may argue against the generic presence of jets in these systems, the large uncertainties prevent us from reaching definitive constraints.
Figure 5: Top-hat jet afterglow light curves at $\nu=6\,{\rm GHz}$ for observers at $\theta_{\rm obs}=0.5\,{\rm rad}$ (solid lines) and on-axis observers (dashed lines) at $z=0.25$, assuming jet Lorentz factor $\Gamma=100$, opening angle $\theta_{j}=0.1\,{\rm rad}$, and the following microphysics parameters: electron equipartition parameter $\epsilon_{e}=0.1$, magnetic equipartition parameter $\epsilon_{B}=10^{-4}$, electron energy power-law index $p=2.2$, and wind velocity $v_{w}=1000\,\,{{\rm km}}\,\,{{\rm s}}^{-1}$. The black circles denote the most constraining upper limits measured for Type I SLSN Gaia2016apd (Coppejans et al., 2018), placing the source at $z=0.25$.
Figure 5 presents the semi-analytic thin-shell approximation solution for the radio synchrotron emission of our model at various assumed progenitor wind mass-loss rates, for both off-axis (solid lines) and on-axis (dashed lines) viewing angles. We assume a characteristic redshift of $z=0.25$ for Type I SLSNe (e.g., Gomez et al., 2024), standard jet properties with $\Gamma=100$ and $\theta_{j}=0.1\,{\rm rad}$, and microphysics parameters (detailed in the caption) inferred from GW170817 afterglow and very long baseline interferometry images (Mooley et al., 2018). If the mass-loss rates are $M_{w}\gtrsim 10^{-5}\,\,{M_{\odot}}\,{\rm yr}^{-1}$, the emission is self-absorbed during the first $\sim 1$ week, leading to an initial rise in the light curves for all observers.
@@
Beyond the general uncertainties in the afterglow and jet parameters, the strong dependence of the flux on the viewing angle (Nakar et al., 2002) and the wind mass-loss rate complicates robust predictions that can support or refute our model. The most stringent upper limits on Type I SLSNe were obtained for Gaia2016apd (Coppejans et al., 2018) at $z=0.1018$ (Yan et al., 2017). The black circles represent these upper limits (adjusted for $z=0.25$), potentially ruling out some scenarios with high mass-loss rates and small viewing angles for this source, but not excluding off-axis observers and lower mass-loss rates within the considered parameters.